# Memory-Augmented Generative AI for Real-time Wireless Prediction in Dynamic Industrial Environments

### v1

@@ Abstract
Accurate and real-time prediction of wireless channel conditions, particularly the Signal-to-Interference-plus-Noise Ratio (SINR), is a foundational requirement for enabling Ultra-Reliable Low-Latency Communication (URLLC) in highly dynamic Industry 4.0 environments. Traditional physics-based or statistical models fail to cope with the spatio-temporal complexities introduced by mobile obstacles and transient interference inherent to smart warehouses. To address this, we introduce Evo-WISVA (Evolutionary Wireless Infrastructure for Smart Warehouse using VAE), a novel synergistic deep learning architecture that functions as a lightweight 2D predictive digital twin of the radio environment. Evo-WISVA integrates a memory-augmented Variational Autoencoder (VAE) featuring an Attention-driven Latent Memory Module (LMM) for robust, context-aware spatial feature extraction, with a Convolutional Long Short-Term Memory (ConvLSTM) network for precise temporal forecasting and sequential refinement. The entire pipeline is optimized end-to-end via a joint loss function, ensuring optimal feature alignment between the generative and predictive components. Rigorous experimental evaluation conducted on a high-fidelity ns-3-generated industrial warehouse dataset demonstrates that Evo-WISVA significantly surpasses state-of-the-art baselines, achieving up to a 47.6% reduction in average reconstruction error. Crucially, the model exhibits exceptional generalization capacity to unseen environments with vastly increased dynamic complexity (up to ten simultaneously moving obstacles) while maintaining amortized computational efficiency essential for real-time deployment. Evo-WISVA establishes a foundational technology for proactive wireless resource management, enabling autonomous optimization and advancing the realization of predictive digital twins in industrial communication networks.
{IEEEkeywords}
Industry 4.0, 5G, 2D Digital Twin, SINR Prediction, GenAI, Deep Learning, Variational Autoencoder (VAE), ConvLSTM, Latent memory module, Spatio-Temporal Modeling.
1 Introduction
[…]
@@ 1 Introduction
Accurate SINR prediction is challenging, particularly in dynamic industrial environments. These settings combine deterministic factors (topology, materials) with highly stochastic elements (mobile obstacles, transient interference, multipath fading). This complexity often yields noisy or incomplete data, making reliable forecasting difficult. Imperfect information leads to inefficient resource use and degraded quality of service. Mission-critical applications require trustworthy forecasting of link reliability, demanding more than simple point predictions [5, 6, 7, 8].
Deep learning (DL) offers a transformative solution, leveraging its ability to learn complex, nonlinear spatio-temporal mappings. While Convolutional Neural Networks (CNNs) capture spatial features and Long Short-Term Memory (LSTM) networks capture temporal dependencies, existing architectures have limitations in dynamic RF environments. CNNs lack memory, and recurrent models often struggle with robustness and uncertainty disentanglement over long prediction horizons [9]. Furthermore, real-time deployment requires efficient models capable of handling continuous data streams and incremental adaptation [10, 1, 4].
These challenges are most acute in Industry 4.0 environments, where mobile obstacles and automated logistics necessitate real-time foresight for Ultra-Reliable Low-Latency Communication (URLLC). To address this, we propose the Evolutionary WISVA (Evo-WISVA) model. This memory-augmented generative architecture is designed for spatio-temporal SINR forecasting, functioning as a lightweight 2D predictive digital twin of the radio environment. Unlike computationally heavy, physics-driven digital twins, Evo-WISVA is a data-driven neural surrogate capable of emulating the wireless environment in real time. By anticipating the evolution of the network state based on continuous measurements and environmental descriptors, Evo-WISVA enables proactive resource management and system-level optimization. This shifts wireless system design from a reactive to a predictive paradigm, advancing the realization of fully autonomous industrial wireless networks.
@@
    End-to-end trainability for wireless digital twins: Unlike many hybrid models, the full pipeline (VAE + Latent Memory Module + ConvLSTM) is trained jointly with a combined loss function, ensuring optimal feature alignment for SINR forecasting.
@@
2 Related Work
Artificial Intelligence (AI) and Machine Learning (ML), particularly Deep Learning (DL), are transforming wireless communication by enabling systems to learn, reason, and optimize complex network functions [11]. The accurate and real-time prediction of wireless channel conditions, specifically Signal-to-Interference-plus-Noise Ratio (SINR) maps, is crucial for next-generation network management and the development of wireless digital twins. This section reviews channel prediction methodologies, highlighting their limitations and how our work, including the proposed Evo-WISVA model, addresses them.
2.1 Traditional Wireless Channel Modeling and Prediction
@@
2.2 Machine Learning for Spatio-Temporal Wireless Channel Prediction
@@
To address these challenges, particularly the need for efficient and automated wireless network design in dynamic industrial settings, our work introduced the Wireless Infrastructure for Smart warehouses using VAE (WISVA) model [28, 29]. WISVA leverages Variational Autoencoders (VAEs) for indoor radio propagation modeling within automated Industry 4.0 warehouses, focusing on 5G wireless bands. As comprehensively detailed in [29], WISVA integrates physics-based tensors (e.g., permittivity, geometry, AP layout) via a three-branch encoder, enabling enhanced modeling of electromagnetic wave interactions. The model is trained on NS-3 simulated SINR heatmaps, learning to generate these heatmaps from latent tensors that encapsulate diverse warehouse configurations. This approach aims to simplify wireless network design by enabling the prediction of SINR heatmaps for both known and previously unseen warehouse layouts, including tasks like denoising and extrapolation, outperforming traditional autoencoders and SDU-Net models [28, 29]. This moves beyond resource-intensive direct measurements or traditional simulations.
Generative models, particularly Variational Autoencoders (VAEs), have shown promise in learning compact latent representations and capturing uncertainty in high-dimensional data [30, 31]. Our earlier WISVA model [29] demonstrated that VAEs can generate SINR heatmaps directly from physics-informed tensors, reducing reliance on costly simulations or exhaustive measurements. However, standard VAEs process data largely frame by frame and fail to capture long-term temporal dependencies, limiting their utility for spatio-temporal forecasting.
To address these gaps, we propose Evo-WISVA, a memory-augmented generative architecture that functions as a 2D predictive digital twin of the wireless environment. Evo-WISVA integrates an attention-driven latent memory module within a VAE–ConvLSTM pipeline, enabling it to retain long-range temporal context while producing robust, denoised, and context-aware spatial features. Unlike static digital twins that merely mirror current conditions, Evo-WISVA anticipates the evolution of SINR maps over time, providing real-time foresight into channel dynamics. This predictive digital twin paradigm is particularly suited for Industry 4.0 scenarios, where wireless reliability must be continuously ensured despite moving obstacles, shifting layouts, and fluctuating interference. In doing so, Evo-WISVA closes the gap between generative spatial modeling and temporal forecasting, establishing a new class of lightweight, map-centric digital twins for wireless networks.
3 Proposed Solution: Synergistic Deep Learning Architecture
@@
Figure 1: Synergistic Deep Learning Architecture for Dynamic Wireless Prediction. The Dynamic_VAE captures memory-augmented spatial representations, which are then fused with raw inputs for temporal refinement by the ConvLSTM, all optimized through a joint loss function.
3.1 Dynamic_VAE Component: Memory-Augmented Spatial Understanding
@@
These modalities are critical as they provide a comprehensive representation of the static and dynamic environmental factors that govern radio frequency (RF) propagation. The Dynamic_VAE processes these inputs through the following stages:
  1. 1.
    Encoder Module: The input $x_{t}$ is fed into the encoder, denoted as $f_{enc}(\cdot)$. This module employs parallel convolutional layers, where each set of layers is dedicated to processing a specific input modality. This parallel processing facilitates modality-specific feature extraction, allowing the model to capture the unique physical influences and characteristics of each map on RF propagation independently. The extracted features are subsequently concatenated and further processed by shared convolutional and fully connected layers. The encoder’s output consists of the mean ($\mu_{t}$) and log-variance ($\log\sigma_{t}^{2}$) of the latent distribution for the current time step $t$:
    $\mu_{t}$ $=f_{enc,\mu}(x_{t})$ (1)
    $\log\sigma_{t}^{2}$ $=f_{enc,\log\sigma}(x_{t})$ (2)
    Here, $f_{enc,\mu}$ and $f_{enc,\log\sigma}$ represent the distinct heads of the encoder network responsible for generating the latent distribution parameters.
  2. 2.
    Reparameterization Trick: To enable backpropagation through the stochastic sampling process, a latent vector $z_{t}$ is sampled from a Gaussian distribution parameterized by $\mu_{t}$ and $\sigma_{t}^{2}$ using the reparameterization trick [32]. This is formulated as:
    $z_{t}=\mu_{t}+\sigma_{t}\odot\epsilon$ (3)
    where $\epsilon\sim\mathcal{N}(0,I)$ is a random sample from a standard normal distribution, and $\odot$ denotes the element-wise product. This operation ensures that the gradients can flow deterministically through the sampling process to optimize the encoder.
  3. 3.
    Latent Memory Module (LMM): The sampled latent vector $z_{t}$ is then passed to the Latent Memory Module. This module is designed to integrate historical context, thereby enhancing the temporal consistency of the VAE’s generative process. The LMM maintains a ‘current_memory_bank‘ (denoted as $\mathcal{M}$) which stores a fixed-size history of past latent states, typically in a First-In-First-Out (FIFO) manner. Crucially, an attention mechanism is employed within the LMM to selectively weigh and combine relevant information from this memory bank with the current latent vector $z_{t}$. The current latent vector $z_{t}$ serves as the query (Q), while past latent states from $\mathcal{M}$ function as both keys (K) and values (V). The attention weights $a_{t,j}$ for each memory entry $m_{j}\in\mathcal{M}$ are computed as:
@@
    where $d_{k}$ is the dimension of the latent vectors (keys). The augmented latent vector $z_{augmented\_t}$ is then derived as a weighted sum of the memory contents:
    $z_{augmented\_t}=z_{t}+\sum_{m_{j}\in\mathcal{M}}a_{t,j}m_{j}$ (5)
    This process ensures that the VAE’s subsequent decoder stage produces reconstructions that are dynamically informed by historical context, thereby mitigating temporal ”flickering” and promoting greater consistency in the predicted sequence. After processing, the memory bank $\mathcal{M}$ is updated with the current latent vector $z_{t}$ for use in the next time step.
  4. 4.
    Decoder Module: The $z_{augmented\_t}$ is then fed into the decoder, denoted as $f_{dec}(\cdot)$. The decoder utilizes a series of deconvolutional (transposed convolutional) layers to progressively upsample the latent representation back to the spatial dimensions of the input. The output of the decoder is the reconstructed SINR map for the current time step, denoted as $\hat{y}_{t}$:
@@
For the entire input sequence, the VAE produces a sequence of reconstructed SINR maps, denoted as vae_pred_seq, with the following dimensions:
$\text{vae\_pred\_seq}\in\mathbb{R}^{\text{Batch}\times\text{Sequence\_Length}\times 1\times H\times W}$ (7)
Additionally, the sequences of means ($\mu_{seq}$) and log-variances ($log\_var_{seq}$) are collected across all time steps for the VAE’s loss calculation.
3.2 ConvLSTM Model Component: Temporal Refinement and Prediction
@@
3.3 Complementary Roles of Latent Memory and ConvLSTM
The synergy between the VAE’s Latent Memory Module (LMM) and the ConvLSTM model is pivotal to the architecture’s effectiveness. While both components address temporal aspects, they operate at distinct levels of abstraction and fulfill complementary roles:
  1. 1.
    Dynamic_VAE Latent Memory Module’s Role: The LMM primarily ensures temporal coherence and contextual understanding within the VAE’s internal generative process. A standard VAE typically processes each input frame independently. When applied to sequences, this frame-by-frame processing can lead to undesirable ”flickering” or inconsistencies in reconstructed outputs, as the VAE lacks explicit knowledge of preceding states. The LMM’s core purpose is to imbue the current frame’s latent representation ($z_{t}$) with relevant historical context from a memory bank of past latent codes. By augmenting $z_{t}$ into $z_{augmented\_t}$, the VAE’s decoder (when generating $\hat{y}_{t}$) is informed not only by the current input $x_{t}$ but also by the evolution of patterns the VAE has previously encoded. This results in a smarter, smoother, and more contextually aware initial estimate of the SINR map sequence (vae_pred_seqvae\_pred\_seq), effectively acting as a high-quality ”prior” or ”initial guess” for the subsequent temporal model.
  2. 2.
    ConvLSTM’s Role: The ConvLSTM, conversely, serves as the primary sequence predictor and refiner. It receives a rich, already pre-processed input (lstm_input_seqlstm\_input\_seq) that judiciously combines both the raw environmental features ($x_{seq}$) and the VAE’s memory-augmented SINR estimate (vae_pred_seqvae\_pred\_seq). The ConvLSTM’s strengths are multifaceted:
    • It excels at learning complex temporal dynamics across the entire sequence of input features and the VAE’s outputs, capturing long-term dependencies and trends.
    • It performs crucial refinement and correction on the VAE’s initial predictions. While the memory-augmented VAE provides an excellent contextualized estimate, it might not capture all subtle, non-linear long-term temporal trends or the full predictive context required for precise forecasting. The ConvLSTM acts as a powerful corrective and predictive layer.
    • It inherently maintains spatial coherence through its convolutional operations within the LSTM cells, which is vital for maintaining the structural integrity of map-based predictions.
    • Ultimately, it generates the final, most accurate spatio-temporal SINR prediction (outputs_seqoutputs\_seq) by leveraging the combined information.
In summary, the VAE with LMM provides a robust, temporally consistent initial representation of the SINR maps. The ConvLSTM then leverages this high-quality starting point (alongside the raw inputs) to learn the precise temporal evolution and produce the final, refined forecast. This division of labor ensures that each component optimally contributes to addressing distinct aspects of the spatio-temporal modeling challenge, rather than performing redundant tasks, leading to enhanced overall prediction performance.
3.4 Joint Loss Calculation and Optimization

### v3

@@ Abstract
Accurate and real-time prediction of wireless channel conditions, particularly the Signal-to-Interference-plus-Noise Ratio, is a foundational requirement for enabling Ultra-Reliable Low-Latency Communication in highly dynamic Industry 4.0 environments. Traditional physics-based or statistical models fail to cope with the spatio-temporal complexities introduced by mobile obstacles and transient interference inherent to smart warehouses. To address this, we introduce Evo-WISVA (Evolutionary Wireless Infrastructure for Smart Warehouse using VAE), a novel synergistic deep learning architecture. This framework ensures real-time foresight into wireless performance, supporting advanced industrial automation and proactive resource management. Evo-WISVA integrates a memory-augmented Variational Autoencoder featuring an Attention-driven Latent Memory Module for robust, context-aware spatial feature extraction, with a Convolutional Long Short-Term Memory network for precise temporal forecasting and sequential refinement. The entire pipeline is optimized end-to-end via a joint loss function, ensuring optimal feature alignment between the generative and predictive components. Rigorous experimental evaluation conducted on a high-fidelity ns-3-generated industrial warehouse dataset demonstrates that Evo-WISVA significantly surpasses state-of-the-art baselines, achieving up to a 47.6% reduction in average reconstruction error. Crucially, the model exhibits exceptional generalization capacity to unseen environments with vastly increased dynamic complexity (up to ten simultaneously moving obstacles) while maintaining amortized computational efficiency essential for real-time deployment. Evo-WISVA establishes a foundational technology for proactive wireless resource management, enabling autonomous optimization and advancing the realization of resilient industrial communication networks.
I Introduction
@@
Accurate SINR prediction is challenging, particularly in dynamic industrial environments [5, 6, 7, 8]. These settings combine deterministic factors (topology, materials) with highly stochastic elements (mobile obstacles, transient interference, multipath fading) [5, 6, 7, 8]. This complexity often yields noisy or incomplete data, making reliable forecasting difficult. Imperfect information leads to inefficient resource use and degraded quality of service. Mission-critical applications require trustworthy forecasting of link reliability, demanding more than simple point predictions [5, 6, 7, 8].
Deep learning (DL) offers a transformative solution, leveraging its ability to learn complex, nonlinear spatio-temporal mappings [9]. While Convolutional Neural Networks (CNNs) capture spatial features and Long Short-Term Memory (LSTM) networks capture temporal dependencies, existing architectures have limitations in dynamic RF environments [9]. CNNs lack memory, and recurrent models often struggle with robustness and uncertainty disentanglement over long prediction horizons [9]. Furthermore, real-time deployment requires efficient models capable of handling continuous data streams and incremental adaptation [10, 1, 4].
These challenges are most acute in Industry 4.0 environments, where mobile obstacles and automated logistics necessitate real-time foresight for Ultra-Reliable Low-Latency Communication (URLLC). To address this, we propose the Evolutionary WISVA (Evo-WISVA) model. This memory-augmented generative architecture is designed for spatio-temporal SINR forecasting, functioning as a neural surrogate capable of emulating the wireless environment in real time. By anticipating the evolution of the network state based on continuous measurements and environmental descriptors, Evo-WISVA enables proactive resource management and system-level optimization. This shifts wireless system design from a reactive to a predictive paradigm, advancing the realization of fully autonomous industrial wireless networks.
@@
    End-to-end trainability for feature alignment: Unlike many hybrid models, the full pipeline (VAE + Latent Memory Module + ConvLSTM) is trained jointly with a combined loss function, ensuring optimal feature alignment between the generative and predictive components for superior SINR forecasting.
@@
II Related Work
Artificial Intelligence (AI) and Machine Learning (ML), particularly Deep Learning (DL), are transforming wireless communication by enabling systems to learn, reason, and optimize complex network functions [11]. The accurate and real-time prediction of wireless channel conditions, specifically Signal-to-Interference-plus-Noise Ratio (SINR) maps, is crucial for next-generation network management and the development of advanced industrial communication networks. This section reviews channel prediction methodologies, highlighting their limitations and how our work, including the proposed Evo-WISVA model, addresses them.
II-A Traditional Wireless Channel Modeling and Prediction
@@
II-B Machine Learning for Spatio-Temporal Wireless Channel Prediction
@@
To address these challenges, particularly the need for efficient and automated wireless network design in dynamic industrial settings, our work introduced the Wireless Infrastructure for Smart warehouses using VAE (WISVA) model [28]. WISVA leverages Variational Autoencoders (VAEs) for indoor radio propagation modeling within automated Industry 4.0 warehouses, focusing on 5G wireless bands. WISVA integrates physics-based tensors (e.g., permittivity, geometry, AP layout) via a three-branch encoder, enabling enhanced modeling of electromagnetic wave interactions. The model is trained on NS-3 simulated SINR heatmaps, learning to generate these heatmaps from latent tensors that encapsulate diverse warehouse configurations. This approach aims to simplify wireless network design by enabling the prediction of SINR heatmaps for both known and previously unseen warehouse layouts, including tasks like denoising and extrapolation, outperforming traditional autoencoders and SDU-Net models [28]. This moves beyond resource-intensive direct measurements or traditional simulations.
Generative models, particularly Variational Autoencoders (VAEs), have shown promise in learning compact latent representations and capturing uncertainty in high-dimensional data [29, 30]. Our earlier WISVA model demonstrated that VAEs can generate SINR heatmaps directly from physics-informed tensors, reducing reliance on costly simulations or exhaustive measurements. However, standard VAEs process data largely frame by frame and fail to capture long-term temporal dependencies, limiting their utility for spatio-temporal forecasting.
To address these gaps, we propose Evo-WISVA, a memory-augmented generative architecture that functions as a data-driven neural surrogate of the wireless environment. Evo-WISVA integrates an attention-driven latent memory module within a VAE–ConvLSTM pipeline, enabling it to retain long-range temporal context while producing robust, denoised, and context-aware spatial features. Unlike models that merely mirror current conditions, Evo-WISVA anticipates the evolution of SINR maps over time, providing real-time foresight into channel dynamics. This capability is particularly suited for Industry 4.0 scenarios, where wireless reliability must be continuously ensured despite moving obstacles, shifting layouts, and fluctuating interference. In doing so, Evo-WISVA closes the gap between generative spatial modeling and temporal forecasting, establishing a new class of lightweight, map-centric prediction tools for proactive network management.
III Proposed Solution: Synergistic Deep Learning Architecture
@@
Figure 1: Synergistic Deep Learning Architecture for Dynamic Wireless Prediction.
III-A Dynamic_VAE Component: Memory-Augmented Spatial Understanding
@@
These modalities are critical as they provide a comprehensive representation of the static and dynamic environmental factors that govern radio frequency (RF) propagation. The Dynamic_VAE processes these inputs through the following four stages:
Encoder Module
The input $x_{t}$ is fed into the encoder $f_{enc}(\cdot)$, which employs parallel convolutional layers dedicated to specific input modalities. This parallel processing facilitates modality-specific feature extraction, capturing the unique physical influences of each map on RF propagation. The extracted features are concatenated and processed by shared layers to output the mean ($\mu_{t}$) and log-variance ($\log\sigma_{t}^{2}$) of the latent distribution for the current time step $t$:
$\mu_{t}$ $=f_{enc,\mu}(x_{t}),\quad\log\sigma_{t}^{2}=f_{enc,\log\sigma}(x_{t})$ (1)
This encoding process acts as a nonlinear spectral filter. Under the ELBO objective $\mathcal{L}_{VAE}=\mathbb{E}_{q(z|x)}[\log p(x|z)]-\beta D_{KL}(q(z|x)||p(z))$, the latent bottleneck $z$ identifies a $d$-dimensional manifold $\mathcal{Z}$ spanning the dynamic subspace. By constraining the latent space to a standard Gaussian prior, the VAE filters dominant stationary eigen-directions ($\lambda_{static}$) that would otherwise induce gradient saturation in the temporal layers.
Reparameterization Trick
To enable backpropagation through the stochastic sampling, a latent vector $z_{t}$ is sampled using the reparameterization trick [31]:
$z_{t}=\mu_{t}+\sigma_{t}\odot\epsilon,\quad\epsilon\sim\mathcal{N}(0,I)$ (2)
where $\odot$ denotes the element-wise product. This ensures that gradients flow deterministically to optimize the encoder parameters.
Latent Memory Module (LMM)
The vector $z_{t}$ is passed to the LMM to integrate historical context and enhance temporal consistency. The LMM maintains a memory bank $\mathcal{M}$ of past latent states. An attention mechanism selectively weighs information from $\mathcal{M}$ using $z_{t}$ as the query ($Q$) and past states as keys ($K$) and values ($V$). The weights $a_{t,j}$ for each entry $m_{j}\in\mathcal{M}$ are computed as:
@@
The augmented latent vector $z_{augmented\_t}$ is then derived as a weighted sum: $z_{augmented\_t}=z_{t}+\sum_{m_{j}\in\mathcal{M}}a_{t,j}m_{j}$. This process informs the reconstruction with historical context, mitigating temporal ”flickering” across the sequence.
Decoder Module
Finally, $z_{augmented\_t}$ is fed into the decoder $f_{dec}(\cdot)$, which utilizes deconvolutional layers to upsample the representation back to the spatial dimensions of the input. The output is the reconstructed SINR map for the current time step:
@@
For the entire sequence, the VAE produces a sequence $\text{vae\_pred\_seq}\in\mathbb{R}^{\text{Batch}\times\text{Sequence\_Length}\times 1\times H\times W}$, alongside means ($\mu_{seq}$) and log-variances ($log\_var_{seq}$) required for the joint loss calculation.
III-B ConvLSTM Model Component: Temporal Refinement and Prediction
@@
III-C Theoretical Justification: Spectral Stability and Operator Synergy
To analyze the stability and predictive reliability of the Evo-WISVA architecture, we examine the spectral properties of the transition operator $\mathcal{T}$. This hybrid design explicitly addresses the fundamental trade-off between temporal gradient stability and spatial inductive bias inherent in industrial RF environments.
Proposition 1 (LMM as a Non-Markovian Gradient Shortcut).
Let $H_{t}$ be the hidden state of the ConvLSTM refinement module. In traditional recurrent architectures, the sensitivity of the loss $\mathcal{L}$ to a state $k$ steps in the past is governed by the temporal Jacobian $J_{t}=\frac{\partial H_{t}}{\partial H_{t-1}}$. By the chain rule and the sub-multiplicative property of the spectral norm $\|\cdot\|_{2}$:
$\left\|\frac{\partial\mathcal{L}}{\partial H_{t-k}}\right\|_{2}\leq\left\|\frac{\partial\mathcal{L}}{\partial H_{t}}\right\|_{2}\prod_{i=0}^{k-1}\sigma_{max}(J_{t-i})$ (11)
In high-mobility industrial tracking, $\sigma_{max}(J)<1$ is required for state stability, which leads to vanishing gradients $\lim_{k\to\infty}\|\frac{\partial\mathcal{L}}{\partial H_{t-k}}\|=0$ [32]. The Evo-WISVA Latent Memory Module (LMM) introduces an additive attention operator $\mathcal{A}$ such that the augmented latent state is $z_{augmented\_t}=z_{t}+\sum a_{t,j}m_{j}$. The gradient path for the memory component is:
$\frac{\partial z_{augmented\_t}}{\partial m_{j}}=a_{t,j}I$ (12)
Because this term is independent of the sequence depth $k$, it ensures the effective spectral radius $\rho(J_{eff})\approx 1$. This mathematical shortcut enables long-term temporal credit assignment for periodic industrial patterns that are unreachable by standard recurrent units.
Proposition 2 (ConvLSTM as a Local Differential Operator).
While the LMM is permutation-invariant and lacks inherent spatial topology, the ConvLSTM module defines the transition as a spatial convolution: $H_{t}=\sigma(\mathcal{K}*H_{t-1}+\mathcal{W}*\text{lstm\_input\_seq}_{t}+b)$. This operator satisfies the spatial continuity requirement derived from the transport equation $\nabla\cdot\vec{J}+\frac{\partial\rho}{\partial t}=0$ [12], ensuring that the SINR value at coordinate $(x,y)$ is functionally dependent on its immediate neighborhood $\mathcal{N}(x,y)$. This enforces a physical smoothness constraint on the latent trajectory, modeling the ”flow” of RF shadows as a continuous differential process.
Synergy and URLLC Convergence.
The Evo-WISVA architecture represents a hybrid Integral-Differential operator. The LMM serves as a Global Integral Operator, $h_{global}=\int\alpha(\tau)z(\tau)d\tau$, capturing long-range shadowing context. The ConvLSTM serves as a Local Differential Operator, $\frac{\partial H}{\partial t}\approx f(H,\nabla H)$, ensuring spatial topology and local coherence. Together, they satisfy the composite update:
$H_{final}=\underbrace{\text{ConvLSTM}(H_{t-1},\hat{y}_{t})}_{\text{Spatial Locality Bias}}+\underbrace{\text{LMM}(\mathcal{M})}_{\text{Temporal Gradient Stability}}$ (13)
This dual-path optimization ensures the error $\epsilon$ remains bounded even for $T\gg 1$, satisfying the strict predictive reliability required for URLLC in dynamic warehouse environments [33]. ∎
III-D Joint Loss Calculation and Optimization